OpenStudy (anonymous):

An LCO Flat Screen TV at Best Buy increased in price from $900 to $1,200. What was the percent of increase?

OpenStudy (anonymous):

900 + increase = 1200 increase = 300 increase as a percentage of original price = increase / original = 300 / 900 = 1/3 = 33 1/3%